# Smoothed Graph Contrastive Learning via Seamless Proximity Integration

### v1

@@ 4 Experiments
Appendix C provides comprehensive details of the experiments. The computational cost of the proposed model, in comparison to the baselines, is also provided in Appendix D.2.
@@ 4.1 Node classification
DGI (Veličković et al., 2019) 82.43±1.3 70.96±2.4 83.13±0.6 91.67±0.7 65.07±1.2 77.15±1.1
GRACE (Zhu et al., 2020) 83.16±1.9 68.68±1.3 85.68±0.3 91.43±0.7 81.89±0.6 89.33±1.7
MVGRL (Hassani & Khasahmadi, 2020) 85.44±1.8 72.51±2.9 86.33±0.7 92.91±0.5 85.90±0.6 91.48±0.9
BGRL (Thakoor et al., 2022) 79.12±2.7 63.77±1.7 84.17±0.7 91.6±0.65 84.27±0.5 92.31±0.4
GBT (Bielak et al., 2022) 76.91±1.8 59.46±2.9 85.92±0.5 91.45±0.8 88.11±0.9 91.29±0.7
[…]
SGCL-B 87.50±1.7 73.65±1.3 86.13±0.6 93.15±0.4 87.89±0.7 92.31±1.3
SGCL-D 85.22±0.8 72.66±2.4 86.27±0.5 93.21±0.4 85.47±1.2 91.86±0.7
@@
DGI (Veličković et al., 2019) 67.07±0.5 68.68±0.6 94.11±0.1
GRACE (Zhu et al., 2020) 67.92±0.4 72.10±0.7 94.11±0.2
MVGRL (Hassani & Khasahmadi, 2020) 60.68±0.5 69.90±0.9 93.87±0.3
BGRL (Thakoor et al., 2022) 63.88±0.2 66.23±0.5 92.94±0.3
GBT (Bielak et al., 2022) 69.05±0.3 65.74±0.4 94.07±0.3
[…]
SGCL-B 69.24±0.3 74.33±0.4 93.55±0.2
SGCL-D 69.03±0.4 74.15±0.2 93.19±0.1
Table 3: Comparison of graph classification accuracies of proposed models vs. baselines (mean ± std).
Model IMDB-Binary PTC MUTAG PROTEINS ENZYMES
InfoGraph (Sun et al., 2019) 73.0±0.9 61.7±1.4 89.0±1.1 74.4±0.3 50.2±1.4
GraphCL (You et al., 2020) 71.1±0.4 63.6±1.8 86.8±1.3 74.4±0.5 55.1±1.6
MVGRL (Hassani & Khasahmadi, 2020) 74.2±0.7 62.5±1.7 89.7±1.1 71.5±0.3 48.3±1.2
AD-GCL (Suresh et al., 2021) 71.5±1.0 61.2±1.4 86.8±1.3 75.0±0.5 42.6±1.1
BGRL (Thakoor et al., 2022) 72.8±0.5 57.4±0.9 86.0±1.8 77.4±2.4 50.7±9.0
LaGraph (Xie et al., 2022) 73.7±0.9 60.8±1.1 90.2±1.1 75.2±0.4 40.9±1.7
CGRA (Duan et al., 2023) 75.6±0.5 65.7±1.8 91.1±2.5 76.2±0.6 61.1±0.9
[…]
@@
Node classification is one of the important downstream tasks, employed to reflect the effectiveness of the learned graph representation. In the first experiment, we employ six small and medium-scale benchmark datasets: Cora, Citeseer, Pubmed, CoauthorCS, Computers, and Photo. Table 1 reports the performance of the proposed SGCL model and compares the results with baseline models. To generate mini-batches in this experiment, we utilize a random-walk sampling, as outlined in Appendix B. A summary of the results derived from other mini-batching approaches is reported in Table 8.
The results indicate that our proposed model outperforms the majority of benchmarks, providing validation for the effectiveness of our proposed learning framework. In comparison to other benchmarks, the ”Computers” graph exhibits a notably high average node degree but a lower degree of homophily (see Table 4). Consequently, this reduces the significance of neighboring nodes in the proposed smoothing approaches, leading to performance degradation compared to MVGRL.
For further investigation and to facilitate a comprehensive comparison with the existing state-of-the-art, we replicated this experiment in a full-batch scenario, adhering to the commonly employed data split in self-supervised learning as provided in the Open Graph Benchmark, as described in Appendix D.1 (Table 6). The observed performance verifies the enhanced capacity achieved through the utilization of the geometric structure inherent in graphs, enabling improved exploration of positive and negative pairs within the conventional contrastive learning framework. It is worth noting that SGCL-T, which leverages Laplacian information using the Tubin smoothing approach, provides better and more promising results, even slightly outperforming the other smoothing approaches.
In the second experiment, we perform experiments on three large-scale graphs: ogbn-arxiv, ogbn-products, and ogbn-proteins. In this experiment, the significance of the mini-batch generation step of the proposed framework becomes more prominent since employing full-batch large-scale graphs can impose considerable demands on GPU memory due to the necessity of loading all node embeddings onto the GPU. In this experiment, we employ a random-walk sampling approach to generate mini-batches. The results presented in Table 2, demonstrate that the proposed SGCL model consistently outperforms other contrastive learning methods on large-scale graphs. It’s worth noting that ogbn-products serves as an excellent benchmark for our proposed models due to two key advantages. Firstly, its high homophily rate increases the likelihood of discovering neighboring nodes of positive pairs as new positive pairs, enhancing the performance of the model. Secondly, by using mini-batch graphs, instead of the full-batch graph with numerous connected components, we can effectively move beyond the extremely small components. This approach provides richer neighboring information, resulting in the generating of more efficient augmented graphs that contribute to the improved performance of the contrastive loss framework.
4.2 Graph classification
Graph classification is another important downstream task, employed to reflect the effectiveness of the learned graph representation. In this experiment, we follow the InfoGraph (Sun et al., 2019) setting for graph classification and compare the accuracy with self-supervised state-of-the-art methods. The results reported in Table 3 indicate that, in comparison to the best-performing state-of-the-art methods, the proposed model demonstrates enhanced accuracy for IMDB-BINARY, PROTEINS, and ENZYMES, while maintaining comparable accuracy on other benchmarks. It’s worth mentioning that the accuracies of all models are reported from their respective published papers, except for the BGRL results, which we reproduced under the same experimental setting.
4.3 Impact of integrating node proximity into GCL
Figure 4: Comparisons of the mean geodesic distance between the predicted and ground truth positives.
In this experiment, we investigate how integrating node proximity information into GCL models, as detailed in Section 2.2, enhances the prediction of positive pairs for each anchor node. Since SGCL leverages node proximity information, the majority of negatives closely located to the ground truth positive appear with lower penalties in the loss function. This is in contrast to conventional GCL, where all negatives, including potential false negatives, receive equal penalties. As such it promotes predictions that, similar to conventional GCL, are either precisely located at the ground truth positive or, unlike it, are at least in the geodesic vicinity of the positive. Therefore, in comparison to conventional GCL, we expect SGCL to exhibit a lower geodesic distance between the predicted and ground truth positives for each anchor node.
For practical analysis, we utilize the graphs outlined in (Karimi et al., 2018), comprising 5000 nodes divided into two classes. In each graph, the homophily parameter $h$ determines the likelihood of a node connecting with a neighbor that shares the same label. We generate $10$ different graphs by randomly selecting 500 nodes and adjusting the parameter $h$ within the range of 0 to 0.9, with intervals of 0.1. Figure 4 illustrates the mean geodesic distances between the predicted and ground truth positives in each graph, characterized by homophily parameter $h$. We observe that the mean geodesic distance in all variants of SGCL is consistently lower than the conventional GCL. This difference is particularly pronounced in more challenging graphs characterized by a higher level of heterophily (lower $h$ values). In essence, as $h$ increases, nodes tend to form connections with those of the same label. Consequently, the neighbors of the ground truth positives become inherently similar to themselves, leading to reduced mean geodesic distances in conventional GCL.
@@ 5 Conclusion
6 Impact statements
This paper contributes to the advancement of the field of Machine Learning. While our work holds potential societal consequences, we do not feel that any specific implications need to be highlighted at this time.
@@ A.1 Graph representation learning
In recent years, graph neural networks (GNNs) have made significant progress, by the emergence of a multitude of methods dedicated to enhancing graph representation learning. These methods have been designed to address various aspects of network embeddings, including proximity, structure, attributes, learning paradigms, and scalability (Wu et al., 2021; Bronstein et al., 2017). Among the notable GNN approaches, Graph Convolutional Networks (GCN) (Kipf & Welling, 2017) is one of the foundational GNNs that uses convolutional operations to capture local and global information from neighboring nodes, making them effective for tasks like node classification. To overcome the constraints associated with conventional graph convolutions and their approximations, the Graph Attention Network (GAT) (Veličković et al., 2018) introduces the notion of masked self-attentional layers, thereby enhancing its capacity to capture crucial node relationships. By integrating an autoregressive moving average (ARMA) filter, GNN-ARMA (Bianchi et al., 2021) extends the functionality of GNNs to adeptly capture global graph structures. GWCN, as proposed in (Xu et al., 2019a; Behmanesh et al., 2022), utilizes graph wavelets as spectral bases for convolution. This innovative approach enables the modeling of both local and global structural patterns within graphs. GRAND (Chamberlain et al., 2021) presents an interesting perspective on graph convolution networks (GCNs) by interpreting them as a solution to the heat diffusion equation. TIDE (Behmanesh et al., 2023) introduces an innovative approach to tackle the oversmoothing challenge in the message-passing-based approaches by leveraging the diffusion equation to enable efficient and accurate long-distance communication between nodes in a graph.
However, it’s essential to emphasize that the majority of these methods depend on supervised data, and this can be a significant limitation in real-world applications due to the difficulties associated with acquiring labeled datasets. Several traditional unsupervised graph representation learning methods are designed to learn meaningful representations of nodes in a graph without the need for labeled data or explicit supervision. DeepWalk (Perozzi et al., 2014) employs random walks and skip-gram modeling to capture local graph structure, while node2vec (Grover & Leskovec, 2016) extends this approach with a versatile biased random walk strategy encompassing breadth-first and depth-first exploration. LINE(Tang et al., 2015) focuses on preserving both first-order and second-order proximity information in large-scale networks, and GraphSAGE (Hamilton et al., 2017) combines random walk sampling and aggregation to capture both local and global graph structure. HOPE (Ou et al., 2016) leverages higher-order proximity information to capture structural patterns beyond pairwise node relationships in graphs.
@@ A.2 Graph contrastive learning
Self-supervised learning (SSL) has emerged as a powerful paradigm for mitigating the challenges posed by expensive, limited, and imbalanced labels. It enables deep learning models to train on unlabeled data, reducing the reliance on annotated labels (Xie et al., 2023).
@@
Contrastive Learning (CL) is a popular SSL technique known for its simplicity and strong empirical performance. Its fundamental objective is to create meaningful representations by pushing dissimilar pairs apart and pulling similar pairs closer together. Graph Contrastive Learning (GCL) extends the concept of CL to the domain of graphs. However, dealing with the irregular structure of graph data presents more complex challenges in designing strategies for constructing positive and negative samples compared to CL applied to visual or natural language data (Wu et al., 2023).
@@
Numerous papers have emerged to address the challenges associated with GCL. These papers primarily focus on sharing valuable insights and practical approaches for three key elements of contrastive learning: data augmentation, pretext tasks, and contrastive objective (Wu et al., 2023).
@@
Deep Graph Infomax (DGI) (Veličković et al., 2019) and InfoGraph (Sun et al., 2019) are two fundamental contrastive learning models that train a node encoder by maximizing mutual information between the node representation and the global graph representation. DGI is designed for node representation learning, whereas InfoGraph focuses on graph-level representations.
@@
MVGRL (Hassani & Khasahmadi, 2020) is one of the recent GCL approaches that accomplishes the learning of both node and graph-level representations by considering two matrices, namely adjacency and diffusion, as congruent views of a standard contrastive framework.
@@
The fundamental of the aforementioned GCL approaches is the maximization of local-global mutual information within a framework. However, they all rely on a readout function to generate the global graph embedding which this function can be overly restrictive and may not always be achievable. Moreover, for approaches like DGI (Veličković et al., 2019), there is no guarantee that the resulting graph embedding can effectively capture valuable information from the nodes, as it may not adequately preserve the distinctive features found in node-level embeddings.
@@
Several GCL approaches, including GRACE (Zhu et al., 2020), GraphCL (You et al., 2020), and CSSL (Zeng & Xie, 2021), deviate from the conventional approach of contrasting local-global mutual information. Notably, these methods do not rely on making assumptions about the use of injective readout functions to generate the graph embedding.
@@
The effectiveness of the GCL models depends on comparing each item with many negative points (He et al., 2020). However, relying on these negative examples is problematic, especially for graphs, where defining negative samples in a meaningful manner is particularly difficult.

### v2

@@ 4 Experiments
We train ’2-layer’ GCN encoders using the SGCL framework for $200$ iterations with the Adam optimizer (learning rate $31e-$ 3). In the downstream task, we perform node and graph classification with $l_{2}$-regularized logistic regression, reporting accuracy and standard deviation after 5000 runs. For mini-batch generation, we employ the random-walk approach with a batch size of 2000, a random walk length of 4, and 3 starting root nodes. Appendix D provides comprehensive details of the experiments. A comprehensive computational complexity analysis is also provided in Appendix E.3.
@@ 4.1 Node classification
In the first experiment, we evaluate the SGCL model on six small to medium-scale benchmark datasets: Cora, Citeseer, Pubmed, CoauthorCS, Computers, and Photo. Table 1 presents the performance results in comparison with baseline models. To generate mini-batches in this experiment, we utilize a random-walk sampling, as outlined in Appendix C. A summary of the results derived from other mini-batching approaches is reported in Table 11.
The results indicate that the SGCL model outperforms the state-of-the-art on most benchmarks, validating the effectiveness of our learning framework. On the "Computers" graph, which has a notably high average node degree but lower homophily (Table 4), the influence of neighboring nodes in the smoothing approaches is reduced, leading to performance degradation compared to CGRA.
@@
DGI [10] 82.3±0.6 71.8±0.7 76.8±0.6 92.15±0.63 83.95±0.47 91.61±0.22
GRACE [11] 83.3±0.4 72.1±0.5 73.63±0.20 91.12±0.20 89.53±0.35 92.78±0.45
MVGRL [12] 83.11±0.12 73.3±0.5 84.27±0.04 92.11±0.12 87.52±0.11 91.74±0.07
BGRL [21] 83.77± 0.57 73.07±0.06 84.62±0.35 93.31±0.13 90.34±0.19 93.17±0.3
G-BT [22] 83.63±0.44 72.95±0.17 84.52±0.12 92.95±0.17 88.14±0.33 92.63±0.44
CGRA [23] 83.8±0.4 69.23±1.19 82.8±0.4 92.8±0.5 90.5±0.4 92.4±0.2
GRLC [24] 83.5±0.5 72.6±0.6 82.1±0.4 90.36±0.27 88.54±0.23 92.3±0.5
ProGCL-weight [25] 81.91±0.12 69.24±0.21 84.89±0.04 93.51±0.06 89.28±0.15 93.30±0.09
ProGCL-mix [25] 83.71±0.04 68.38±0.3 84.64±0.03 93.67±0.12 89.55±0.16 93.64±0.13
GraphMAE2 [26] 84.5±.0.6 73.4±0.3 81.4±0.5
AUGCL [27] 88.94±0.44 93.43±0.32
GREET [28] 83.81±0.87 73.08±0.84 80.29±1.00 94.65±0.18 87.94±0.35 92.85±0.31
SGCL-T 84.33±0.45 74.94±0.79 84.25±0.35 92.25±0.15 87.21±0.42 93.12±0.7
SGCL-B 84.78±0.3 74.30±1.4 84.1±0.25 92.33±0.4 89.75±0.8 93.72±0.12
SGCL-D 84.17±0.43 75.72±0.59 85.12±0.3 92.14±0.26 86.11±0.3 92.87±0.6
The observed performance verifies the enhanced capacity achieved through the utilization of the geometric structure inherent in graphs, enabling improved exploration of positive and negative pairs within the conventional contrastive learning framework. More evaluation on heterophilic graphs can be found in Appendix E.2.
Furthermore, we evaluate the proposed framework on three large-scale graphs: ogbn-arxiv, ogbn-products, and ogbn-proteins. Here, the importance of the mini-batch generation step becomes more prominent, as full-batch processing of large-scale graphs can impose considerable demands on GPU memory by requiring all node embeddings to be loaded onto the GPU. We employ a random-walk sampling approach to generate mini-batches.
@@
DGI [10] 67.07±0.5 68.68±0.6 94.11±0.1
GRACE [11] 67.92±0.4 72.10±0.7 94.11±0.2
MVGRL [12] 60.68±0.5 69.90±0.9 93.87±0.3
BGRL [21] 63.88±0.2 66.23±0.5 92.94±0.3
GBT [22] 69.05±0.3 65.74±0.4 94.07±0.3
GraphMAE2 [26] 68.95±0.4 74.32±0.5
SGCL-T 70.89±0.2 75.97±0.1 94.64±0.2
SGCL-B 70.34±0.4 74.33±0.4 93.55±0.2
SGCL-D 70.52±0.3 74.15±0.2 93.19±0.1
The results presented in Table 2, demonstrate that the SGCL consistently outperforms other contrastive learning methods on large-scale graphs. Rssults for GraphMAE2 have been reproduced using the standard data split to ensure fair comparisons. It’s worth noting that ogbn-products serves as a valuable benchmark for our proposed models providing two key advantages. Firstly, its high homophily rate increases the likelihood of identifying neighboring nodes of positive pairs as new positive pairs, thereby enhancing the performance of the model. Secondly, by using mini-batch graphs instead of the full-batch graph with numerous connected components, we can effectively bypass the extremely small components. This approach offers richer neighboring information, leading to the generation of more effective augmented graphs and enhancing the performance of the contrastive loss framework.
4.2 Graph classification
Graph classification is another important downstream task, employed to reflect the effectiveness of the learned graph representation. In this experiment, we follow the InfoGraph [29] setting for graph classification and compare the accuracy with self-supervised state-of-the-art methods. The results reported in Table 3 indicate that, in comparison to the best-performing state-of-the-art methods, the proposed model demonstrates enhanced accuracy for IMDB-BINARY, PROTEINS, and ENZYMES, while maintaining comparable accuracy on other benchmarks. It’s worth mentioning that the accuracies of all models are reported from their respective published papers, except for the BGRL results, which we reproduced under the same experimental setting.
Table 3: Comparison of graph classification accuracies of proposed models vs. baselines.
Model IMDB-Binary PTC-MR MUTAG PROTEINS ENZYMES
InfoGraph [29] 73.0±0.9 61.7±1.4 89.0±1.1 74.4±0.3 50.2±1.4
GraphCL [30] 71.1±0.4 63.6±1.8 86.8±1.3 74.4±0.5 55.1±1.6
MVGRL [12] 74.2±0.7 62.5±1.7 89.7±1.1 71.5±0.3 48.3±1.2
AD-GCL [31] 71.5±1.0 61.2±1.4 86.8±1.3 75.0±0.5 42.6±1.1
BGRL [21] 72.8±0.5 57.4±0.9 86.0±1.8 77.4±2.4 50.7±9.0
LaGraph [32] 73.7±0.9 60.8±1.1 90.2±1.1 75.2±0.4 40.9±1.7
ProGCL-mix [25] 71.6±0.6 88.7±1.4 74.5±0.4
CGRA [23] 75.6±0.5 65.7±1.8 91.1±2.5 76.2±0.6 61.1±0.9
AUGCL [27] 72.4±0.8 89.2±1.0 75.7±0.4
SGCL-T 75.2±2.8 64.0±1.6 89.0±2.3 79.4±1.9 65.3±3.6
@@ 5 Conclusion
Acknowledgements
The authors acknowledge the anonymous reviewers for their valuable suggestions and Johannes Lutzeyer for insightful discussions. Parts of this work were supported by the ERC Consolidator Grant 101087347 (VEGA) and the ANR AI Chair AIGRETTE.
@@ A.1 Graph representation learning
In recent years, graph neural networks (GNNs) have made significant progress, by the emergence of a multitude of methods dedicated to enhancing graph representation learning. These methods have been designed to address various aspects of network embeddings, including proximity, structure, attributes, learning paradigms, and scalability [33, 34]. Among the notable GNN approaches, Graph Convolutional Networks (GCN) [2] is one of the foundational GNNs that uses convolutional operations to capture local and global information from neighboring nodes, making them effective for tasks like node classification. To overcome the constraints associated with conventional graph convolutions and their approximations, the Graph Attention Network (GAT) [4] introduces the notion of masked self-attentional layers, thereby enhancing its capacity to capture crucial node relationships. By integrating an autoregressive moving average (ARMA) filter, GNN-ARMA [35] extends the functionality of GNNs to adeptly capture global graph structures. GWCN, as proposed in [36, 37], utilizes graph wavelets as spectral bases for convolution. This innovative approach enables the modeling of both local and global structural patterns within graphs. GRAND [38] presents an interesting perspective on graph convolution networks (GCNs) by interpreting them as a solution to the heat diffusion equation. TIDE [5] introduces an innovative approach to tackle the oversmoothing challenge in the message-passing-based approaches by leveraging the diffusion equation to enable efficient and accurate long-distance communication between nodes in a graph.
However, it’s essential to emphasize that the majority of these methods depend on supervised data, and this can be a significant limitation in real-world applications due to the difficulties associated with acquiring labeled datasets. Several traditional unsupervised graph representation learning methods are designed to learn meaningful representations of nodes in a graph without the need for labeled data or explicit supervision. DeepWalk [39] employs random walks and skip-gram modeling to capture local graph structure, while node2vec [40] extends this approach with a versatile biased random walk strategy encompassing breadth-first and depth-first exploration. LINE[41] focuses on preserving both first-order and second-order proximity information in large-scale networks, and GraphSAGE [42] combines random walk sampling and aggregation to capture both local and global graph structure. HOPE [43] leverages higher-order proximity information to capture structural patterns beyond pairwise node relationships in graphs.
@@ A.2 Graph contrastive learning
Self-supervised learning (SSL) has emerged as a powerful paradigm for mitigating the challenges posed by expensive, limited, and imbalanced labels. It enables deep learning models to train on unlabeled data, reducing the reliance on annotated labels [8].
@@
Contrastive Learning (CL) is a popular SSL technique known for its simplicity and strong empirical performance. Its fundamental objective is to create meaningful representations by pushing dissimilar pairs apart and pulling similar pairs closer together. Graph Contrastive Learning (GCL) extends the concept of CL to the domain of graphs. However, dealing with the irregular structure of graph data presents more complex challenges in designing strategies for constructing positive and negative samples compared to CL applied to visual or natural language data [7].
@@
Numerous papers have emerged to address the challenges associated with GCL. These papers primarily focus on sharing valuable insights and practical approaches for three key elements of contrastive learning: data augmentation, pretext tasks, and contrastive objective [7].
@@
Deep Graph Infomax (DGI) [10] and InfoGraph [29] are two fundamental contrastive learning models that train a node encoder by maximizing mutual information between the node representation and the global graph representation. DGI is designed for node representation learning, whereas InfoGraph focuses on graph-level representations.
@@
MVGRL [12] is one of the recent GCL approaches that accomplishes the learning of both node and graph-level representations by considering two matrices, namely adjacency and diffusion, as congruent views of a standard contrastive framework.
@@
The fundamental of the aforementioned GCL approaches is the maximization of local-global mutual information within a framework. However, they all rely on a readout function to generate the global graph embedding which this function can be overly restrictive and may not always be achievable. Moreover, for approaches like DGI [10], there is no guarantee that the resulting graph embedding can effectively capture valuable information from the nodes, as it may not adequately preserve the distinctive features found in node-level embeddings.
@@
Several GCL approaches, including GRACE [11], GraphCL [30], and CSSL [44], deviate from the conventional approach of contrasting local-global mutual information. Notably, these methods do not rely on making assumptions about the use of injective readout functions to generate the graph embedding.
@@
The effectiveness of the GCL models depends on comparing each item with many negative points [45]. However, relying on these negative examples is problematic, especially for graphs, where defining negative samples in a meaningful manner is particularly difficult.